# Training-free Camera Control for Video Generation

### v2

@@ 4.1 Experimental Settings
The major results presented in this paper are grounded on zeng2023PixelDance . To ensure a fair comparison with other state-of-the-art methods, we employ SVD as base model for quantitative evaluation. Inversion and generation steps among all methods are set to 25. Our method doesn’t require any additional training utilizing camera trajectories.
@@
We compare CamTrol with state-of-the-art works: AnimateDiff guo2023animatediff , MotionCtrl wang2023motionctrl and CameraCtrl he2024cameractrl , all three methods are finetune on SVD using specific data. In quantitative evaluation, FVD unterthiner2018FVD , FID heusel2017FID and IS saito2020IS are used to assess video generation quality, while CLIPSIM wu2021CLIPSIM quantifies the similarity between generated video and input prompt. With regard to the accuracy of camera motion, we adopt ParticleSFMzhao2022particlesfm and produce estimated camera trajectories from generated videos, with the use of Absolute Trajectory Error(ATE) measuring their differences compared to ground truth. Relative Pose Error(RPE) is calculated to assess between consecutive frames how well the relative motions match expected ones including their transition(RPE-T) and rotation part(RPE-R). Settings of evaluation dataset follow those established in MotionCtrl wang2023motionctrl , but include more complex trajectories. Specifically, we extract 51 distinct trajectories from RealEstate10krealestate10k , each paired with 10 prompts mentioned in MotionCtrl wang2023motionctrl , resulting in 510 samples in total for assessment. As AnimateDiff guo2023animatediff lacks the ability to handle complicated trajectories, it is only included in qualitative analysis part evaluating with 8 basic camera motions. Quantitative comparisons contain AnimateDiff guo2023animatediff can be found in Appendix A.3.
@@ 4.2 Comparisons with State-of-the-art Methods
Quantitative Evaluation
Quantitative evaluations are shown in Table  1. Building upon the same base model as MotionCtrl wang2023motionctrl and CameraCtrl he2024cameractrl , our method demonstrates superior performance across all quantitative metrics concerning both video quality and camera motion accuracy. Despite being training-free, CamTrol outperforms those that rely heavily on extensive finetuning and large-scale annotated data, attaining the lowest score in ATE, RPE-T and RPE-R. Benefit from explicit camera movements modeling and motion inversion, our method is capable to produce videos align best with those tricky trajectories while maintaining their visual qualities.
Table 1: Quantitative comparisons.
Method Video Quality Motion Accuracy
MotionCtrl wang2023motionctrl 3576.40 239.10 7.58 0.2933 4.006 1.086 0.106
CameraCtrl he2024cameractrl 2922.99 243.98 8.07 0.2915 4.200 1.487 0.080
SVD+CamTrol 2832.59 227.36 8.07 0.3100 3.917 0.947 0.017
Figure 3: Qualitative comparisons with finetuned methods. CamTrol generates videos that adhere to desired camera movements, without compromising their dynamics and rationality.
Qualitative Analysis
Qualitative comparisons of different methods are illustrated in Fig. 3. All methods generate videos align well with desired camera motion, however, the constraints imposed by camera trajectories are sometimes too strong, that output videos have to compromise their quality and rationality to satisfy these movements. Some examples that demonstrate this are results of Zoom Out and Roll Anticlockwise, which looks more like flat images being zoomed or rolled on 2D plane instead of videos recorded in 3D spaces. The reason might lies on the limitations of proprietary dataset used for finetuning realestate10k ; yu2023mvimgnet , where most scenes are static, making pretrained model forget about knowledge of dynamic contents. In contrast, CamTrol makes full use of the powerful dynamics in base model and generate plausible videos with camera perspective change utilizing layout prior guidance, handles these circumstances gracefully. More comparisons at Appendix A.3.
4.3 Ablation Study
@@
To demonstrate that changes of camera motion are attributed to our method rather than the innate capability of video model, we conduct ablation study to assess its effectiveness. We add prompts describing certain camera moves(e.g. zooms out), letting video model understand by itself. The results are shown in Fig. 4. It could be observed that even provided with prompts indicating how camera should move, base model fail to produce correct results. Instead, CamTrol  is able to implement designated motion control without any instructions from text prompts.
@@
$t_{0}$ is a crucial factor that influences the trade-off between generated video’s diversity and its faithfulness to camera motions requirements. To investigate its effect on the output, we conduct experiments with various $t_{0}$ values, relevant results are shown in Fig. 7. As illustrated, videos generated with larger $t_{0}$ tend to conform better to camera motion requirements, but suffer from decrease in dynamics; On the contrary, smaller $t_{0}$ leads to more plausible generations but fails to meet camera’s requirements, as latents at these timesteps carry more randomness.
Generalization to Diverse Base Models
Our proposed CamTrol can be seamlessly plugged and played with most pretrained video diffusion models to achieve training-free camera control. We present visual results showcasing its applications based on different video base models, including SVD SVD and VideoFusion luo2023videofusion , in Fig. 8. Our approach remains effective applied to alternative video base models, demonstrating its strong robustness and generalization ability. More results with other base models can be found in Appendix A.5.
@@ 4.4 Further Applications
Figure 6: Effect of $t_{0}$. Smaller $t_{0}$ encourages
dynamics while larger $t_{0}$ preserves camera
movements (Pedestal Down).
Figure 7: Camera control at different scales. CamTrol supports camera movements over various scales.
Figure 8: Applied onto SVD SVD and VideoFusion luo2023videofusion . CamTrol could be plugged and played with most video diffusion models to offer camera movement control.
@@
One of the most advantages of our method is it can generate videos rotating around some objects and produce outputs similar to 3D generation models sv3d ; im3d . While these 3D models need large-scale training on 3D dataset and could only handle inputs in specific styles, our approach is able to deal with any type of images and achieve this in a completely zero-shot manner. Some results are shown in 1. We offer more results on multi-view synthesis in Appendix A.2.
@@
By combining different basic camera trajectories, CamTrol can support hybrid camera movements and endow generated video with cinematic charm. Besides this, explicit motion modeling also equips CamTrol with the abilities to support trajectories containing precise coordinates, which means it can generate videos presenting any complicated camera movement. Results about hybrid and complex motions are shown in Fig. 1 and Appendix A.2.
@@
CamTrol supports camera movements at controllable scales. We provide some results in Fig. 7. By specifying different magnitudes of camera’s extrinsic matrix within point cloud spaces, rendered images will exhibit varying degrees of motion, leading to videos with distinct scales of camera movements. This further demonstrate the powerful controllability of CamTrol, and provides a new pathway for video’s customized camera control.
@@ 5 Conclusion
In this paper, we propose a training-free and robust method CamTrol to offer camera control for off-the-shelf video diffusion models. It consists of two-stage procedure including explicit camera motion modeling in 3D point cloud space and video generation utilizing layout prior of noisy latents. Compared to previours work, CamTrol  does not require any additional finetuning on camera-annotated datasets, or self-supervised training via data augmentation, instead, it could be plugged and played with most video diffusion models to generate camera controllable videos with single image or text prompt as input. Comprehensive experimental results demonstrate the effectiveness and robustness of CamTrol for controlling camera motion for videos. Besides, we show that CamTrol produces impressive results in generating videos with 3D rotations, complicated trajectories, and different moving scales.
Appendix A Appendix
A.1 Definitions of Basic Camera Motions
@@
Table 2: Definitions of basic camera motions.
@@
Hybrid Arbitrary Combination of other motions.
A.2 More Results on Camera Control
In this section, we showcase additional qualitative results of CamTrol on various camera movements. Fig. 9 shows video frames with camera move Rotate Anticlockwise and Rotate Clockwise. These outputs share sort of similarity with outputs of 3D generation models, as they all exhibit in a turning-table like way, which camera rotates around some objects. The difference here is that 3D model, as only trained on specific datasets, could only generate outputs in certain styles, e.g., single static object with no background. Instead, our model could handle arbitrary image as input, and generate a rotating-around video with proper dynamics. From this aspect, our method could be seen as a infinite source of attaining 3D data. And by utilizing our method with stronger backbones, video foundation models could truly become the largest source of 3D data as it should be.
Besides, we provide additional results on hybrid motions and basic motions, which is illustrated in Fig. 10, Fig. 11 and Fig. 13, respectively. We also present pre-loaded complex camera motions in Fig. 12.
A.3 More Comparisons to State-of-the-Art Methods
More comparison results with state-of-the-art methods are shown in Fig. 14 and Fig. 15. Compared to other works, our method matches well with input texts and camera requirements, meanwhile has wide-ranging diversities on generated contents.
We also present quantitative comparisons that include AnimateDiff guo2023animatediff as part of the evaluation. As AnimateDiff guo2023animatediff lacks the ability to handle complicated trajectories, experiments are conducted focusing on 8 basic trajectories described in A.1. To expand the diversity of the models evaluated, we employ CamTrol on zeng2023PixelDance for assessment. With regard to basic trajectories, we apply UCF-101 soomro2012UCF as the reference of calculating FID and FVD, from where we randomly sample 125 prompts per camera motion type and form 1000 samples in total for assessment. Relevant results are exhibited in Table 3.
Table 3: Quantitative comparisons on basic trajectories.
Method Video Quality
FVD $\downarrow$ FID $\downarrow$ IS $\uparrow$ CLIP-SIM $\uparrow$
AnimateDiff guo2023animatediff 1582.93 73.79 13.39 0.3196
MotionCtrl wang2023motionctrl 1768.25 78.26 14.60 0.3143
CameraCtrl he2024cameractrl 1383.68 76.18 14.04 0.3123
CamTrol+zeng2023PixelDance 1134.71 69.70 15.87 0.3253
A.4 Ablation on Depth Optimization
Ablation results relevant to the effectiveness of depth optimization are presented in Table 4. While it affects little on CLIPSIM, depth optimization exhibits benefits on other video quality assessments(FVD, FID, IS) and camera trajectory criteria(ATE, RPE). Experiments are conducted on complicated 10 trajectories extracted from RealEstate10k realestate10k , each with 10 prompts mentioned in MotionCtrl wang2023motionctrl . We assume that depth optimization offers better alignment between adjacent views rendered from point cloud, which benefits both perceptual quality and camera motion accuracy of generated videos.
Table 4: Ablation on depth coefficient optimization.
[…]
w/o optimization 5033.44 219.01 5.67 0.2949 4.093 1.131 0.046
w/ optimization 4977.88 218.36 5.72 0.2947 3.906 0.978 0.049
A.5 More Results on other Base Video Models
Our methods could be plugged and played with arbitrary video diffusion models to achieve camera control. We showcase its performance based on two alternative video generation backbones including SVD SVD and VideoFusion luo2023videofusion , relevant results are exhibited in Fig. 16.
Figure 9: Results with camera move Rotate Anticlockwise and Rotate Clockwise. Our method could generated 3D rotation-like videos with dynamics.
Figure 10: Results with camera move Hybrid I. Here we combine Zoom Out, Truck Left, Pedestal Up, Pan Right and Tilt Down together, form a cinematic like effect on generated videos.
Figure 11: Results with camera move Hybrid II. Here the hybrid motion is defined as Zoom In first, then Pedestal Up.
Figure 12: Results with complex camera motions.
Figure 13: Results with basic camera motions.
Figure 14: Comparison to state-of-the-art methods on Truck Left, Truck Right, Pedestal Up, Pedestal Down.
Figure 15: Comparison to state-of-the-art methods on Zoom In, Zoom Out, Roll Anticlockwise, Roll Clockwise.
Figure 16: Results with other video generation base models including SVD SVD and VideoFusion luo2023videofusion . Our method works well with these backbones.

### v3

@@ 4 Experiments
Table 1: Quantitative comparisons. Our method attains comparable performance with finetuned methods in both video generation quality and camera motion alignment.
Method Video Quality Motion Accuracy
SVD 1107.93 68.51 7.21 0.3095 4.23 1.79 0.021
MotionCtrl+SVD 810.59 69.03 7.17 0.3076 4.19 1.07 0.012
CameraCtrl+SVD 951.80 67.59 6.82 0.3138 4.22 1.17 0.013
CamTrol+SVD 778.46 68.06 7.05 0.3110 4.17 1.07 0.010
Reference - - - - 3.60 0.89 0.008
Figure 3: Qualitative comparisons with finetuned methods. CamTrol’s outputs align well with complex trajectories from reference video, while others fail to perceive subtle camera pose changes. We provide videos in the supplementary materials for a clearer comparison.
Table 2: Computational analysis of inference process, evaluated under unified settings.
SVD MotionCtrl CameraCtrl CamTrol ($t_{0}=10$)
Max GPU memory(MB) 11542 31702 26208 11542
Time (s) pre-process - - - 56
inference 11 32 42 8
Figure 4: Generalization comparisons. CamTrol can avoid domain collapse that rises from overfitting on certain datasets, adapting to more general scenes (Left, where input does not resemble real scenes), and maintain video’s dynamics while adhering to desired camera movements (Right).
@@ 4.1 Experimental Settings
We compare our method with state-of-the-art works including MotionCtrl (Wang et al., 2023) and CameraCtrl (He et al., 2024). To ensure a fair comparison, we employ SVD (Blattmann et al., 2023a) as base model for all methods. SVD is originally trained on resolution of $576\times 1024$, but SVD-based CameraCtrl only support $320\times 576$. Since changing the original resolution leads to suboptimal generation quality, we use $576\times 1024$ for MotionCtrl and CamTrol, then resize their outputs into $320\times 576$ for calculating metrics. For all methods, the number of frames and the decoding size of SVD are set to 14. We use 25 steps for both inversion and generation processes.
@@
In quantitative evaluation, FVD (Unterthiner et al., 2018), FID (Heusel et al., 2017) and IS (Saito et al., 2020) are used to assess video generation quality, and CLIPSIM (Wu et al., 2021) quantifies the similarity between generated video and input prompt. For the accuracy of camera motion, we adopt ParticleSFM(Zhao et al., 2022) and produce estimated camera trajectories from generated videos, with the use of Absolute Trajectory Error(ATE) measuring their differences compared to ground truth. Relative Pose Error (RPE) is calculated to assess between consecutive frames how well the relative motions match expected ones including their transition (RPE-T) and rotation part (RPE-R). Specifically, we randomly sample 500 prompt-trajectory pairs from RealEstate10k (Zhou et al., 2018), and use the corresponding videos as reference for calculating FVD and FID. Since SVD is an image-to-video model, we generate the first frames using Stable Diffusion (Rombach et al., 2022) by text prompts. We also compare the results produces by vanilla SVD as a reference. For camera motion accuracy, we provide the evaluations on ground truth videos as a lower bound of these metrics.
@@ 4.2 Comparisons with State-of-the-art Methods
Quantitative Evaluation
Quantitative evaluations are shown in Table  1. In the table, the best results are in bold, and the second best is underlined. The performance of vanilla SVD (without motion control) is indicated as SVD, while the lower bound of motion metrics provided by ground truth videos is indicated as Reference. In terms of video quality, CamTrol attains comparable performance with methods finetuned on the RealEstate10k dataset. For motion accuracy, CamTrol also gets lowset score in ATE and RPE-T/R. The quantitative evaluation demonstrates CamTrol’s ability to generate videos with both accurate camera motion and high visual quality.
Qualitative Analysis
Qualitative comparisons are illustrated in Fig. 3. The reference trajectory includes zoom, pan and roll. While MotionCtrl and CameraCtrl fail to perceive subtle camera motions, resulting in simple pan movement, CamTrol is able to follow the complex trajectory and generate video with correct move. We also evaluate the generalization ability of different methods in generating more general scenes and dynamic content, the results are shown in Fig. 4. Since both MotionCtrl and CameraCtrl are finetuned on limited scenes (i.e., real estate videos) with static content, they can hardly generalize to other scenes, such as videos in non-realistic style or videos that include people. As illustrate in Fig. 4, their motion controls in both situations are misaligned with input movements. Furthermore, finetuning on such datasets leads to the loss of dynamics, which is a crucial element in video generation. In comparison, CamTrol preserves most of the prior knowledge from video base models, enabling it to handle general scenes as well as generate dynamic content. Relevant videos are available in the supplementary materials.
Computational Analysis
We provide the computational analysis in Table 2, including the maximum GPU memory required and the time consumption for all methods during inference process. The evaluations are conducted under unified settings. We test on $576\times 320$, with 25 generation steps, the number of frames and decoding size of SVD are set to 14. As a training-free method, CamTrol requires no extra GPU memory during inference process compared with base model. This makes it save 10-20GB of GPU memory than other methods under the same circumstances, and makes it applicable running on a single RTX 3090. The major consumption of CamTrol comes from the time rendering multi-view images. Since this process is currently sequential in time, i.e., $t_{0}\to t_{1}\to t_{2}$, a more parallel approach may be more time-efficient. We will leave this to future works. The results of $576\times 1024$ resolution and more detailed settings can be found in the Appendix.
4.3 Ablation Study
Table 3: Quantitative effect of $t_{0}$.
$t_{0}$ Video Quality Motion Accuracy
$t_{0}=5$ 1079.88 68.52 7.14 0.3100 4.17 1.09 0.012
$t_{0}=10$ 778.46 68.06 7.05 0.3110 4.17 1.07 0.010
$t_{0}=15$ 754.14 67.98 7.00 0.3107 4.13 1.02 0.008
Figure 7: Effect of $t_{0}$. Smaller $t_{0}$ encourages
dynamics while larger $t_{0}$ preserves camera
movements (Pedestal Down).
Figure 8: 3D rotation videos at different motion scales. CamTrol supports camera movements over various scales.
@@ 4.3 Ablation Study
To demonstrate that changes of camera motion are attributed to our method rather than the innate capability of video model, we conduct ablation study to assess its effectiveness. We add prompts describing certain camera moves (e.g. zooms out), letting video model understand by itself. The results are shown in Fig. 5. It could be observed that even provided with prompts indicating how camera should move, base model fail to produce correct results. Instead, CamTrol  is able to implement designated motion control without any instructions from text prompts. In Table 1, the comparison with vanilla SVD also demonstrate CamTrol’s effectiveness.
@@
$t_{0}$ is a crucial factor that influences the trade-off between generated video’s diversity and its faithfulness to camera motions requirements. To investigate its effect on the output, we conduct experiments with various $t_{0}$ values, relevant results are shown in Fig. 8. As illustrated, videos generated with larger $t_{0}$ tend to conform better to camera motion requirements, but suffer from decrease in dynamics; On the contrary, smaller $t_{0}$ leads to more plausible generations but fails to meet camera’s requirements, as latents at these timesteps carry more randomness. We also provide quantitative evaluations of different $t_{0}$ in Table 3. As $t_{0}$ increase, CamTrol produces videos alike to static, camera-moving scenes (which have lower FVD as we use RealEstate10k as reference videos) with higher accuracy in motion control.
Generalization to Diverse Situations
Our proposed CamTrol can be seamlessly plugged and played under various scenarios, accommodating different base models, resolutions, generating length, all in training-free style. We present visual results of its applications under different settings, including CogVideoX (Yang et al., 2024b)(diffusion transformer model, $720\times 480$ resolution, 49 frames) and VideoFusion (Luo et al., 2023)(decomposed diffusion process, $128\times 128$ resolution, 16 frames), in Fig. 10. Our approach remains effective applied to alternative video base models, resolutions and generating lengths, demonstrating its strong robustness and generalization ability.
@@ 4.4 Further Applications
Figure 9: Multi-trajectory video generation. CamTrol is able to generate videos with different trajectories for one scene.
Figure 10: Applied onto CogVideoX (Yang et al., 2024b) and VideoFusion (Luo et al., 2023). CamTrol can be plugged and played under various situations, accommodating different base models, different resolutions, different generating lengths, all in a training-free manner.
@@
One of the most advantages of our method is it can generate videos with rotating moves and produce outputs similar to 3D generation models (Voleti et al., 2024; Melas-Kyriazi et al., 2024). While these 3D models need large-scale training on 3D dataset and could only handle inputs in specific styles, our approach is able to deal with any type of images and achieve this in a completely zero-shot manner. One example of this is shown in Fig. 8.
@@
By combining different basic camera trajectories, CamTrol can support hybrid camera movements and endow generated video with cinematic charm. Besides this, explicit motion modeling also equips CamTrol with the ability to support trajectories containing precise extrinsics, and generate videos presenting any complicated camera movement. Videos avaiable at https://lifedecoder.github.io/CamTrol/.
Multi-trajectory Video Generation
One natural application of our method is to generate multi-trajectory videos for one scene. Since the point cloud is fixed once the reconstruction process finished, the consistency of the content between different trajectories is guaranteed, and we can easily produce multiple camera-moving videos of this scene. We present this application in Fig. 9. More results on multi-trajectory video generation could be found in the supplementary materials.
@@
CamTrol supports camera movements at controllable scales. By specifying different magnitudes of camera’s extrinsic matrix within point cloud spaces, rendered images will exhibit varying degrees of motion, leading to videos with distinct scales of camera movements. This further demonstrate the powerful controllablity of CamTrol, and provides a new pathway for video’s customized camera control. We provide relevant visualization in Fig. 8.
@@ 5 Conclusion
In this paper, we propose a training-free and robust method CamTrol to offer camera control for off-the-shelf video diffusion models. It consists of two-stage procedure including explicit camera motion modeling in 3D point cloud space and video generation utilizing layout prior of noisy latents. Compared to previous work, CamTrol  does not require any additional finetuning on camera-annotated datasets or self-supervised training via data augmentation, simultaneously, comprehensive experiments demonstrate its superior performance in both generation video quality and motion alignment against other state-of-the-art methods. Furthermore, we show the ability of CamTrol  generalizing to various scenarios, as well as its impressive applications including unsupervised generation of 3D rotation video, scalable motion control and dealing with complicated camera trajectories.
Appendix A More Results on Camera Control
We showcase additional qualitative results of CamTrol on various camera movements. As pdf format is not the best way to exhibit videos, we put them on our demo page:
If not accessible, please refer to the supplementary materials for a quicker view to the demo page.
This demo page includes CamTrol-generated videos including basic camera motions (including Zoom, Tilt, Pan, Pedestal, Truck, Roll, Rotate, Hybrid, Complicated, detailed definitions are in B), hybrid motions (Zoom In first, then Pedestal Up, Zoom Out + Pedestal Up + Truck Left + Tilt Down + Pan Right) and complicated motions generated from precise camera extrinsics (extracted from RealEstate10k (Zhou et al., 2018)).
Besides, it contains 3D rotation videos generated unsupervisedly from video base models (both objects and scenes). These outputs share sort of similarity with outputs of 3D generation models, as they all exhibit in a turning-table like way, which camera rotates around some objects. The difference here is that 3D model, as trained on specific datasets, could only generate outputs in certain styles, e.g., single static object with no background. Instead, our model could handle arbitrary image as input, and generate a rotating-around video with proper dynamics. From this aspect, our method could be seen as a infinite source of attaining 3D data. And by utilizing our method with stronger backbones, video foundation models could truly become the largest source of 3D data as it should be.
Furthermore, it contains additional results mentioned in ablation studies, e.g. controlling camera motions at different scales, and the effectiveness of using layout prior compared with the raw output given by video base model.
Appendix B Definitions of Basic Camera Motions
@@
To get consistent images from multiple views, we set camera motion as a trajectory of extrinsic matrix $\left\{\mathbf{P}_{1},...,\mathbf{P}_{N-1}\right\}$, each including a rotation matrix and translation matrix representing camera’s pose and position. For hybrid motions, CamTrol supports both spatial (i.e. moving several basic moves simultaneously) and temporal (concatenation of basic motions sequentially) combinations. In the case of complicated trajectories, one can directly use precise parameters for camera’s extrinsic matrix as input to control video’s motion. Additionally, if these parameters not available, a reference video with corresponding move can also serve as input. With the help of SFM algorithms (e.g. COLMAP  [^1] https://github.com/colmap/colmap, ParticleSFM (Zhao et al., 2022)), camera motion can be easily estimated and used for imitation. In this sense, CamTrol can also be seen as an unsupervised method to achieve video motion transfer.
Appendix C Visual Comparisons with State-of-the-arts
Sec. 3 showcases some qualitative comparisons between CamTrol and different state-of-the-art approaches. For better visualization and comparison, we provide more video results in the supplementary materials. Please check them if needed.
Appendix D Implementation Details
For text prompt input, we use Stable Diffusion v2-1  [^2] https://huggingface.co/stabilityai/stable-diffusion-2-1 or Stable Diffusion XL  [^3] https://huggingface.co/stabilityai/stable-diffusion-xl-base-1.0 to generate the initial image. The inpainting model we apply is Stable Diffusion inpainting model proposed by runway  [^4] https://huggingface.co/runwayml/stable-diffusion-inpainting, and backward step of inpainting is set to 25. We deploy ZeoDepth  [^5] https://github.com/isl-org/ZoeDepth as depth estimation model. The classifier-free guidance scale follows default setting of base models themselves. We use SVD’s default setting of 6 fps for video generation, and process reference videos to 6 fps for FVD and FID calculation. The complete procedure of CamTrol is elucidated in Algorithm 1.
For computational analysis, we set both number of frames and decoding size to 14, generation steps to 25. We do not apply xformers in all approaches. The computational analysis at $576\times 1024$ resolution is shown in Table 4. SVD-based CameraCtrl only support resolution at $320\times 576$.
Table 4: Computational analysis of on $\textbf{576}\times\textbf{1024}$.
SVD MotionCtrl CameraCtrl CamTrol ($t_{0}=10$)
Max GPU memory(MB) 34236 71096 - 34236
Time (s) pre-process - - - 149
inference 32 54 - 22
Table 5: Definitions of camera motions. We follow the terminology in cinematography to describe different camera movements.
@@ Appendix D Implementation Details
Hybrid Combinations Spatial and temporal combination of basic motions.
Complicated Arbitrary Camera extrinsic matrix or a reference video as input.
Appendix E Choice of 3D Representation
Input: Text prompt $p$, camera motion $\mathbf{P}$, input image $\mathbf{I}_{0}$ (optional).
// Stage I: Camera Motion Modeling:
1 for i=1,…,N-1 do
2       $\tilde{\mathbf{I}}_{i}=\text{inpainting}(\mathbf{I}_{i-1},\mathbf{P}_{i},p)$ ;
3       $\tilde{\mathbf{D}}_{i}=\text{depth}(\tilde{\mathbf{I}}_{i})$ ;
4       while not converged do
5             $d_{i}=\operatorname*{argmin}_{d}\left(\sum_{M}\left\|\phi([\tilde{\mathbf{I}}_{i},d\tilde{\mathbf{D}}_{i}],\mathbf{K},\mathbf{P}_{i})-\mathcal{P}_{i-1}\right\|\right)$
6       end while
7      $\mathcal{P}_{i}=\phi([\tilde{\mathbf{I}}_{i},d_{i}\tilde{\mathbf{D}}_{i}],\mathbf{K},\mathbf{P}_{i})$ ;
8       $\mathbf{I}_{i}=\psi(\mathcal{P}_{i},\mathbf{K},\mathbf{P}_{i})$ ;
9
10 end for
// Stage II: Layout Prior Generation:
11 $\mathbf{V}_{0}\leftarrow\left\{\mathbf{I}_{i}\right\}_{i=0}^{N-1}$ ;
12 Sample random noise $\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ ;
13 Motion inversion $\mathbf{V}_{t_{0}}\leftarrow\sqrt{\bar{\alpha}_{t_{0}}}\mathbf{V}_{0}+\sqrt{1-\bar{\alpha}_{t_{0}}}\epsilon$ ;
14 for t=$t_{0}$,…,1 do
15       $\mathbf{V}_{t-1}\leftarrow\sqrt{\alpha_{t-1}}\left(\frac{\mathbf{V}_{t}-\sqrt{1-\alpha_{t}}\epsilon^{(t)}_{\theta}(\mathbf{V}_{t})}{\sqrt{\alpha_{t}}}\right)+\sqrt{1-\alpha_{t-1}-\sigma^{2}_{t}}\epsilon^{(t)}_{\theta}(\mathbf{V}_{t})+\sigma_{t}\epsilon$
16 end for
Algorithm 1 Training-free camera control for video generation
In Sec. 3.1, we designate point cloud as the intermediate 3D representation for explicit camera motion modeling. Doubts may arise why we do not use more complex 3D representation which might be more useful. Here we take the most prevalent 3D representation: 3D Gaussian Splatting  [^6] https://repo-sam.inria.fr/fungraph/3d-gaussian-splatting/, as example to elaborate the benefit of using point cloud in three aspects:
Firstly, concerning the input, point cloud is able to construct the whole scene from one single input image combining techniques of depth estimation and inpainting. GS, though also an explicit 3D representation, requires optimization following images from different views, which means it is neither capable of handling single input image, nor can it leverages 2D inpainting to complete the scene gradually.
Secondly, from the aspect of time, point cloud can directly lift 2D points onto 3D spaces, while 3DGS demands optimization on each scenario. As a training-free method, our method takes nearly no time to generate a video after multi-view images are acquired, but would need more time if 3DGS were applied.
Lastly, from the task itself, the target of stage I is not precisely reconstruct the 3D scene but only offer a rough layout guidance, in this context, rendered images from point cloud qualify enough and no further refinement on 3D reconstruction is necessary.
From the analysis, point cloud is quite qualified serving as a rough layout guidance in a relatively quick speed, without any further optimization and redundant multi-view images as input. The use of point cloud allows our algorithm to be totally training-free and optimization-free, simultaneously being able to produce camera-moving videos quickly with merely one single image or text prompt as input.
Appendix F Details about Generating 3D Rotation Videos
3D generation models (Voleti et al., 2024; Melas-Kyriazi et al., 2024; Shi et al., 2023; Chen et al., 2024; Han et al., 2024) are trained on highly-regulated 3D datasets, these datasets are hard to collect, and consist only a limited variety of data types (e.g. single static objects with no background). As a consequence, 3D generation models exhibit very constrained output distribution and could only produce results in certain styles. CamTrol avoid these problems of arduous data collecting, laborious finetuning and output collapse by utilizing the layout prior hidden in video foundation models. Not only does it require no training, but this advantage also benefits CamTrol from inheriting most of the prior knowledge inside video foundation models, e.g., the diversity of scenarios, the dynamic of moving objects, temporal consistency, etc. Thus CamTrol is able to generate dynamic 3D content, both objects and scenes, in a totally unsupervised and training-free manner, this is what other method cannot achieve yet.
Compared with regulated datasets, the problem of processing wild pictures in 3D is that some of the parameters is unknown. In Sec. 3.1, we’ve mentioned that the camera intrinsic matrix $\mathbf{K}$ and initial extrinsic matrix $\mathbf{P}_{0}$ are set by convention as they’re usually intractable. Another crucial parameter concerning 3D video generation is the distance between camera and the content of input image (denote as $f$), note that input image could be synthetic or real. Considering most camera rotations are done around the center point, we extract a patch from the very center of the input image, conducting depth estimations on the it and define the distance $f$ as the averaged depth. The rotation and transition matrix during camera moving can be formed as :
$\mathcal{R}_{y}$ $=\begin{bmatrix}\cos\theta_{i}&0&-\sin\theta_{i}\\ 0&1&0\\ \sin\theta_{i}&0&\cos\theta_{i}\end{bmatrix},\quad t=\begin{bmatrix}f\sin\theta_{i}\\ 0\\ f-f\cos\theta_{i}\end{bmatrix},$ (6)
$\text{where}\quad f=\frac{1}{|P|}\sum_{(j,k)\in P}D(x_{0}+j,y_{0}+k).$
Here $i\in[0,N-1]$ and $\theta_{i}$ refers to rotation angle around $y$ axis at step $i$, $D$ denotes depth estimation of image, $P$ represents the patch around central point $(x_{0},y_{0})$. In our experiment, we choose $(j,k)\in[-10,10]$ as the size of patch.
Appendix G Motion Blur, Problems and Solutions
Videos produced by CamTrol need to satisfy certain camera movement, and some drastic perspective changes might cause visible trails, recognised as motion blur of objects or scenes. This phenomenon will appear to be more indispensable when video base model holds a relatively small generation length (e.g. 16 frames) as well as the motion scale becomes larger (e.g. tilt up for 60 degrees or more). To avoid blur issues in controlling video camera motion, we propose several solutions as below:
  1. 1.
    According to the analysis above, the blur issue is caused by limited generation frames and large camera movement, thus the most intuitive solution is to either cut down motion scale or utilize a more capable generation model. For severe perspective changes, the optimal approach would involve employ video foundation models that support larger generation length (e.g. CogVideoX (Yang et al., 2024b) supports generating video with 49 frames). This allows model to manage motions of equivalent magnitude while experiencing a smaller moving range between adjacent frames, thereby brings effective alleviation to blur problems.
  2. 2.
    One can also stack the results of multiple generations to form a complete outcome, i.e., treating the last frame of previous generation as the starting frame for the next, and integrate them as a whole. This approach is more suitable when utilizing a image-to-video (I2V) base model. Since most open-source video foundation models are text-to-video (T2V), one may consider increasing the step of camera motion inversion $t_{0}$, which guarantees more fidelity towards input images’ content (and motion).
  3. 3.
    Besides the above two approaches, it is as well a common and convenient choice to apply frame interpolation towards the output. Lots of off-the-shelf frame interpolation models are open-sourced and could be found on github.
  4. 4.
    Lastly, if video length is not being able to alter, it may be necessary to increase the fps of generated videos for better visualization. Although larger fps leads to shorter video duration, it simultaneously makes the visual persistence brought by motions less pronounced, which reduce blur visually.
In our experiments, we take raw output in all settings.